OpenStudy (anonymous):

Find an equation for the line tangent to the graph of y=tan x at the point (pi/4, 1). b) Find an equation for the line tangent to the graph of y=arc tan x at the point 9=(1, pi/4)

OpenStudy (anonymous):

derivatives of tan x is sec^2(0)

OpenStudy (anonymous):

slope = sec^2(1)

OpenStudy (kira_yamato):

Part 1

OpenStudy (kira_yamato):

Part 2

OpenStudy (anonymous):

How do you know that sec^2(pi/4)=2?

OpenStudy (kira_yamato):

sec x = 1/(cos x) So, sec^2 x = 1/cos^2 x cos π/4 = 1/√2 cos^2 π/4 = 1/2 sec^2 π/4 = 2